OpenStudy (anonymous):

A right triangle has a hypotenuse of length 39 and a leg of length 15. What is the length of the other leg? If necessary round your answer to two decimal places.

OpenStudy (marissalovescats):

Hi and welcome to OpenStudy! Are you familiar with the Pythagorean Theorem?

OpenStudy (marissalovescats):

Pythagorean states that a^2+b^2=c^2. Where a and b are the 2 shorter sides of the triangle and c is the hypotenuse. This only works with right triangles, it helps us find missing sides. So here they gave us a smaller leg and a hypotenuse so we have: 15^2+b^2=39^2 So first we want to know this: 15^2 = 225 and 39^2=1521 so now what we have is this: 225+b^2=1521 and now we solve for b: We subtract 225 from each side to get: b^2=1296 and now we take the square root of both sides to get b: b=36. So our missing leg is 36.